# Quantitative estimates for singularity for conjugate equations driven by linear fractional transformations

### v3

@@ Abstract.
We consider the conjugate equation driven by two families of finite maps on the unit interval satisfying a compatibility condition. This framework contains de Rham’s functional equations. We give sufficient conditions for singularity of the solution with quantitative estimates in the case that the equation is driven by non-affine maps, in particular, linear fractional transformations.
@@ 1. Introduction
Let $N\geq 2$ and let $\mathcal{I}_{N}:=\{0,1,\cdots,N-1\}$. Let $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ be two families of $N$ maps on $[0,1]$ satisfying a compatibility condition. We consider the following conjugate equation for $\varphi:[0,1]\to[0,1]$:
$g_{i}\circ\varphi=\varphi\circ f_{i},\ i\in\mathcal{I}_{N}.$ (1.1)
See Section 2 below for the definition of the compatibility condition. An alternative expression is given by
$\varphi(x)=\begin{cases}g_{0}(\varphi(f_{0}^{-1}(x)))&0\leq x\leq f_{0}(1)\\ \vdots&\\ g_{N-1}(\varphi(f_{N-1}^{-1}(x)))&f_{N-1}(0)\leq x\leq 1\end{cases}.$ (1.2)
@@
Assume that $N=2$, $10<p<$ 1, $f_{0}(x)=x/2$, $f_{1}(x)=(x+1)/2$, $g_{0}(x)=px$ and $g_{1}(x)=(1-p)x+p$. Then, (1.2) is expressed as
@@
There exists a unique continuous strictly increasing solution $\varphi_{p}$ for this equation, and $\varphi_{p}$ is singular if $p\neq 1/2$ and $\varphi_{p}(x)=x$ if $p=1/2$. $\varphi_{p}$ is called the Lebesgue singular function. This function is also a solution of the functional equation by Matkowski and Wesołowski [16, 24, 25].
Minkowski’s question-mark function [23] is also obtained by a solution of (1.1) for the case that $(f_{0}(x),f_{1}(x),g_{0}(x),g_{1}(x))=\left(x/(x+1),1/(2-x),x/2,(x+1)/2\right)$. It can also be expressed in the form
@@
The main purpose of this paper is to investigate some properties of the solution of (1.1). Specifically, we give sufficient conditions for the solution to be singular. Since $\varphi$ is an increasing function on $[0,1]$, we deal with the probability measure $\mu_{\varphi}$, which is the Borel probability measure on $[0,1]$ whose distribution function is $\varphi$. We consider the case that both of $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are non-affine. This case is more difficult to analyze than the case that both $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are affine.
It is interesting to know a necessary and sufficient condition of $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ for the singularity. This topic has been considered in [2, 3, 10, 13, 18, 25]. In [27], the author dealt with the case that $\{g_{i}\}_{i}$ are linear fractional transformations and $\{f_{i}\}_{i}$ are affine. Our results, Theorems 3.2 and 3.3 below, deal with the case that $\{f_{i}\}_{i}$ are not necessarily affine. We do not assume the differentiability of $\{f_{i}\}_{i}$. For the proof, we use a martingale approach as in [27], but here we need more delicate quantitative estimates.
The structure of this paper is as follows. In Section 2, we give some terminology and basic results of the conjugate equation on the unit interval. In Section 3, we deal with the singularity. Finally, we give three numerical examples in Section 4.
@@ 2.1. Existence and uniqueness
We say that a family of maps $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ on $[0,1]$ is a compatible system or satisfies a compatibility condition if they are strictly increasing continuous functions such that $h_{0}(0)=0$, $h_{i-1}(1)=h_{i}(0)$, $11\leq i\leq N-$ 1, $h_{N-1}(1)=1$, and, a compatible system $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ is a D-system on $[0,1]$ if
$\mathcal{D}_{h}:=\{h_{i_{1}}\circ\cdots\circ h_{i_{n}}(j):i_{1},\cdots i_{n}\in\mathcal{I}_{N},j\in\{0,1\}\}$
@@
$x=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1).$ (2.1)
If $x\in\mathcal{D}_{h}\setminus\{0,1\}$, then, there exist two such infinite sequences. Otherwise, there exists exactly one such infinite sequence.
@@
$\lim_{n\to\infty}|h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)-h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)|=0.$ (2.2)
Assume this fails. By the definition of $\{h_{i}\}_{i}$, $\left(h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)\right)_{n}$ is increasing and $\left(h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\right)_{n}$ is decreasing. Since $\mathcal{D}_{h}$ is dense in $[0,1]$, there exists $k_{1},\cdots,k_{N}$ and $j\in\{0,1\}$ such that
$h_{i_{1}}\circ\cdots\circ h_{i_{N}}(0)\leq\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)<h_{k_{1}}\circ\cdots\circ h_{k_{N}}(j)$
$<\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\leq h_{i_{1}}\circ\cdots\circ h_{i_{N}}(1).$
This cannot occur. Hence (2.2) holds.
Case 1. Assume $x\in\mathcal{D}_{h}\setminus\{0,1\}$. Assume that $x=h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)$. Since $h_{0}(0)=0$ and $x>0$, we can assume that $i_{n}\geq 1$. Let $i_{k}:=0$ for every $k\geq n+1$. Then, (2.1) holds. By replacing $i_{n}$ with $i_{n}-1$ and let $i_{k}:=N-1$ for every $k\geq n+1$, then, (2.1) holds. Assume that $x=h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)$. Since $h_{N-1}(1)=1$ and $x<1$, we can assume that $i_{n}\leq N-2$. Let $i_{k}:=N-1$ for every $k\geq n+1$. Then, (2.1) holds. By replacing $i_{n}$ with $i_{n}+1$ and let $i_{k}:=0$ for every $k\geq n+1$, then, (2.1) holds.
@@
Case 3. Assume $x\notin\mathcal{D}_{h}$. Then, for each $n\geq 1$, there exists a unique finite sequence $(i_{n,1},\cdots,i_{n,n})$ such that
@@
Then, $i_{n-1,k}=i_{n,k}$ for every $k\leq n-1$. Let $i_{k}:=i_{n,k}$. By (2.2), (2.1) holds. If (2.1) holds, then, for every $N\geq 1$,
$h_{i_{1}}\circ\cdots\circ h_{i_{N}}(0)\leq\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)=x=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\leq h_{i_{1}}\circ\cdots\circ h_{i_{N}}(1).$
@@
$h_{i_{1}}\circ\cdots\circ h_{i_{N}}(0)<x<h_{i_{1}}\circ\cdots\circ h_{i_{N}}(1).$
@@
We first see that
$\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(j)\right)=g_{i_{1}}\circ\cdots\circ g_{i_{n}}(j),\ i_{1},\cdots,i_{n}\in\mathcal{I}_{N},\ j\in\{0,1\},$ (2.3)
by induction in $n$. Hence, $\varphi(\mathcal{D}_{f})=\mathcal{D}_{g}$.
Step 1. We define $\varphi(x)$ for $x\notin\mathcal{D}_{f}$. By the assumption, there exists an infinite sequence $(i_{n})_{n}$ of $\mathcal{I}_{N}$ such that $x\in(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0),f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1))$. Since $\left[g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0),g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)\right]$ is decreasing with respect to $n$,
@@
$\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0)=\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1).$ (2.4)
We let $\varphi(x):=\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0)$.
Step 2. We show $\varphi$ is strictly increasing on $[0,1]$. By (2.3), $\varphi$ is strictly increasing on $\mathcal{D}_{f}$. Let $x<y$. By (2.4), for some $n$,
@@
Step 3. We show that $\varphi$ is continuous on $[0,1]$. Assume that there is a discontinuous point $x_{0}\in(0,1)$ of $\varphi$. Then,
$b_{0}:=\lim_{x\to x_{0}+0}\varphi(x)>\lim_{x\to x_{0}-0}\varphi(x)=:a_{0}.$
By Step 2, $\varphi([0,1])\cap(a_{0},b_{0})=\emptyset$, in particular, $\mathcal{D}_{g}\cap(a_{0},b_{0})=\emptyset$. This contradicts the density of $\mathcal{D}_{g}$ in $[0,1]$. The same argument is applied to the case that $x_{0}\in\{0,1\}$. ∎
It is natural to consider the inverse function of the solution $\varphi$. We easily see that
@@
Assume that $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are D-systems on $[0,1]$. Let $\varphi$ be a unique solution of (1.1). Then, the inverse function $\varphi^{-1}$ of $\varphi$ is well-defined and satisfies the following conjugate equation
@@
which is equivalent with (1.1) where $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are exchanged.
By [4, Proposition 3.1], if $\varphi$ is singular, then, $\varphi^{-1}$ is also singular.

### v4

@@ Abstract.
We consider the conjugate equation driven by two families of finite maps on the unit interval satisfying a compatibility condition. This framework contains de Rham’s functional equations. We give sufficient conditions for singularity of the solution with quantitative estimates in the case where the equation is driven by a family of non-affine maps and a family of linear fractional transformations.
@@ 1. Introduction
Let $N\geq 2$ and let $\mathcal{I}_{N}\coloneqq\{0,1,\ldots,N-1\}$. Let $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ be two families of $N$ maps on $[0,1]$ satisfying a compatibility condition. We consider the following conjugate equation for $\varphi\colon[0,1]\to[0,1]$:
(1.1) $g_{i}\circ\varphi=\varphi\circ f_{i},\ i\in\mathcal{I}_{N}.$
See the beginning of Section 2 below for the definition of the compatibility condition. An alternative expression is given by
(1.2) $\varphi(x)=\begin{cases}g_{0}(\varphi(f_{0}^{-1}(x)))&0=f_{0}(0)\leq x\leq f_{0}(1)\\ \vdots&\\ g_{N-1}(\varphi(f_{N-1}^{-1}(x)))&f_{N-1}(0)\leq x\leq f_{N-1}(1)=1\end{cases}.$
@@
Assume that $N=2$, $0<p<1$, $f_{0}(x)=x/2$, $f_{1}(x)=(x+1)/2$, $g_{0}(x)=px$ and $g_{1}(x)=(1-p)x+p$. Then, (1.2) is expressed as
@@
There exists a unique continuous strictly increasing solution $\varphi_{p}$ for this equation, and the function $\varphi_{p}$ is singular if $p\neq 1/2$, and $\varphi_{p}(x)=x$ if $p=1/2$. See [34, Subsection 3.4] for the proof. The function $\varphi_{p}$ is often called the Lebesgue singular function. This function is sometimes referred to by the name of Riesz and Nágy ([1, 3]), Salem ([8]), or Takács ([3]). This function is also a solution of the functional equation by Matkowski and Wesołowski [24, 25].
Minkowski’s question-mark function was initially considered in [23]. This function can be obtained by a solution of (1.1) for the case where $f_{0}(x)=x/(x+1)$, $f_{1}(x)=1/(2-x)$, $g_{0}(x)=x/2$, and $g_{1}(x)=(x+1)/2$. It can also be expressed in the form
@@
The main purpose of this paper is to investigate sufficient conditions for the solution of (1.1) to be singular. Necessary and sufficient conditions of $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ for the singularity of the solution have been considered in [4, 5, 12, 15, 18, 25]. We consider the case where both $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are non-affine maps. This case is more difficult to analyze than the case where both $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are affine maps. In [27], the author dealt with the case where $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are linear fractional transformations and $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ are affine maps. Our results, Theorems 3.2 and 3.3 below, deal with the case where $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ are not necessarily affine maps. We do not assume the differentiability of $\{f_{i}\}_{i\in\mathcal{I}_{N}}$. Since $\varphi$ is an increasing function on $[0,1]$, we deal with the probability measure $\mu_{\varphi}$, which is the Borel probability measure on $[0,1]$ whose distribution function is $\varphi$. We give sufficient conditions for $\dim_{H}\mu_{\varphi}<1$, which implies the singularity for $\varphi$. For the proof, we use an approach using martingales as in [27], but here we need more delicate quantitative estimates for $\dim_{H}\mu_{\varphi}$.
The structure of this paper is as follows. In Section 2, we give some terminology and basic results of the conjugate equation on the unit interval. In Section 3, we give our main results and their proofs. Finally, we give three numerical examples in Section 4.
@@ 2.1. Existence and uniqueness
We say that a family of maps $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ on $[0,1]$ is a compatible system or that it satisfies a compatibility condition if they are strictly increasing continuous functions such that $h_{0}(0)=0$, $h_{i-1}(1)=h_{i}(0)$, $1\leq i\leq N-1$, $h_{N-1}(1)=1$, and that a compatible system $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ is a D-system on $[0,1]$ if
$\mathcal{D}_{h}\coloneqq\left\{h_{i_{1}}\circ\cdots\circ h_{i_{n}}(j)\colon i_{1},\ldots,i_{n}\in\mathcal{I}_{N},j\in\{0,1\}\right\}$
@@
(2.1) $x=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1).$
If $x\in\mathcal{D}_{h}\setminus\{0,1\}$, then there exist two such infinite sequences. Otherwise, there exists exactly one such infinite sequence.
@@
(2.2) $\lim_{n\to\infty}|h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)-h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)|=0.$
Assume this fails. By the definition of $\{h_{i}\}_{i\in\mathcal{I}_{N}}$, the sequence $\left(h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)\right)_{n}$ is increasing and $\left(h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\right)_{n}$ is decreasing. Since $\mathcal{D}_{h}$ is dense in $[0,1]$, there exist $M\in\mathbb{N}$ and $k_{1},\ldots,k_{M}$ as well as $j\in\{0,1\}$ such that
$h_{i_{1}}\circ\cdots\circ h_{i_{M}}(0)\leq\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)<h_{k_{1}}\circ\cdots\circ h_{k_{M}}(j)$
$\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\leq h_{i_{1}}\circ\cdots\circ h_{i_{M}}(1).$
This cannot occur since the sequence $\left(h_{i_{1}}\circ\cdots\circ h_{i_{n}}(j)\right)_{i_{1},\dots,i_{n},j}$ is in lexicographic order, more precisely, $h_{i_{1}}\circ\cdots\circ h_{i_{n}}(j)\leq h_{i_{1}^{\prime}}\circ\cdots\circ h_{i_{n}^{\prime}}(j^{\prime})$ if and only if $\displaystyle j+\sum_{\ell=1}^{n}i_{\ell}N^{n-\ell}\leq j^{\prime}+\sum_{\ell=1}^{n}i_{\ell}^{\prime}N^{n-\ell}$. Hence (2.2) holds.
Case 1. Assume $x\in\mathcal{D}_{h}\setminus\{0,1\}$. Assume that $x=h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)$. Since $h_{0}(0)=0$ and $x>0$, we can assume that $i_{n}\geq 1$. Let $i_{k}\coloneqq 0$ for every $k\geq n+1$. Then, the first equality of (2.1) holds. By this and (2.2), the second equality of (2.1) holds. Thus (2.1) holds. By replacing $i_{n}$ with $i_{n}-1$ and letting $i_{k}\coloneqq N-1$ for every $k\geq n+1$, the equations in (2.1) hold. Assume that $x=h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)$. Since $h_{N-1}(1)=1$ and $x<1$, we can assume that $i_{n}\leq N-2$. Let $i_{k}\coloneqq N-1$ for every $k\geq n+1$. Then, (2.1) holds.
It remains valid if we replace $i_{n}$ with $i_{n}+1$ and let $i_{k}\coloneqq 0$ for every $k\geq n+1$. There are no more solutions of the infinite sequences satisfying (2.1) since the sequence $\left(h_{i_{1}}\circ\cdots\circ h_{i_{n}}(j)\right)_{i_{1},\dots,i_{n},j}$ is in lexicographic order.
@@
Case 3. Assume $x\notin\mathcal{D}_{h}$. Then, for each $n\geq 1$, there exists a unique finite sequence $(i_{n,1},\ldots,i_{n,n})$ such that
@@
Then, $i_{n-1,k}=i_{n,k}$ for every $k\leq n-1$. Let $i_{k}\coloneqq i_{n,k}$. By (2.2), the equations in (2.1) hold. If (2.1) holds, then for every $M\geq 1$,
$h_{i_{1}}\circ\cdots\circ h_{i_{M}}(0)\leq\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)=x=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\leq h_{i_{1}}\circ\cdots\circ h_{i_{M}}(1).$
@@
$h_{i_{1}}\circ\cdots\circ h_{i_{M}}(0)<x<h_{i_{1}}\circ\cdots\circ h_{i_{M}}(1).$
@@
First we show the existence.
Step 1. We define $\varphi(x)$ for $x\in\mathcal{D}_{f}$. Let
(2.3) $\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(j)\right)\coloneqq g_{i_{1}}\circ\cdots\circ g_{i_{n}}(j),\ i_{1},\ldots,i_{n}\in\mathcal{I}_{N},\ j\in\{0,1\}.$
This is well-defined since $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are compatible systems. Hence, $\varphi(\mathcal{D}_{f})=\mathcal{D}_{g}$.
Step 2. We define $\varphi(x)$ for $x\notin\mathcal{D}_{f}$. By the assumption and Lemma 2.1, there exists a unique infinite sequence $(i_{n})_{n}$ of $\mathcal{I}_{N}$ such that $x\in(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0),f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1))$. Since $\left[g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0),g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)\right]$ is decreasing with respect to $n$,
@@
$\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0)=\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1).$
We let $\displaystyle\varphi(x)\coloneqq\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0)$.
Step 3. We show that $\varphi$ is strictly increasing on $[0,1]$. By (2.3), $\varphi$ is strictly increasing on $\mathcal{D}_{f}$. Let $x<y$. By the density of $\mathcal{D}_{f}$ in $[0,1]$, we can apply (2.2), and obtain that
$\lim_{n\to\infty}f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0)=\lim_{n\to\infty}f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1).$
Hence it holds that for some $n$ and $i_{1},\dots,i_{n},j_{1},\dots,j_{n}\in\mathcal{I}_{N}$,
@@
Step 4. We show that $\varphi$ is continuous on $[0,1]$. Assume that there is a discontinuous point $x_{0}\in(0,1)$ of $\varphi$. Then,
$b_{0}\coloneqq\lim_{x\to x_{0}+0}\varphi(x)>\lim_{x\to x_{0}-0}\varphi(x)\eqqcolon a_{0}.$
By Step 3, $\varphi([0,1])\cap(a_{0},b_{0})=\emptyset$, in particular, $\mathcal{D}_{g}\cap(a_{0},b_{0})=\emptyset$. This contradicts the density of $\mathcal{D}_{g}$ in $[0,1]$. The same argument is applied to the case where $x_{0}\in\{0,1\}$.
Second we show the uniqueness.
Step 1. We show that if $\varphi$ is a solution of (1.1), then $\varphi(j)=j,j\in\{0,1\}$. By (1.1) and the assumption that $f_{0}(0)=0$, $\varphi(0)=\varphi(f_{0}(0))=g_{0}(\varphi(0))$. Denote the $n$-fold composition of $g_{0}$ by $g_{0}^{n}$. Then, by induction in $n$, we see that $\varphi(0)=g_{0}^{n}(\varphi(0))$ for every $n\geq 1$. Since $g_{0}$ is an increasing function on $[0,1]$, $g_{0}^{n}$ is also an increasing function on $[0,1]$ for each $n$. Therefore, $0=g_{0}^{n}(0)\leq g_{0}^{n}(\varphi(0))\leq g_{0}^{n}(1)$. By this and (2.2),
$\varphi(0)=\lim_{n\to\infty}g_{0}^{n}(\varphi(0))=\lim_{n\to\infty}g_{0}^{n}(1)=0.$
In the same manner, we can show that $\varphi(1)=1$.
Step 2. Since $\varphi(j)=j,j\in\{0,1\}$, we see that
(2.4) $\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(j)\right)=g_{i_{1}}\circ\cdots\circ g_{i_{n}}(j),\ i_{1},\ldots,i_{n}\in\mathcal{I}_{N},\ j\in\{0,1\},$
by induction in $n$.
Since $\mathcal{D}_{f}$ is dense in $[0,1]$ and $\varphi$ is assumed to be continuous, we obtain the uniqueness. ∎
We give an example of a compatible system $\{h_{0},h_{1}\}$ which is not a $D$-system. This is suggested by an anonymous referee.
Example 2.4.
Let
$h_{0}(x)\coloneqq\begin{cases}\frac{1}{2}x&0\leq x\leq\frac{1}{4},\\ 2x-\frac{3}{8}&\frac{1}{4}\leq x\leq\frac{1}{2},\\ \frac{3}{8}x+\frac{7}{16}&\frac{1}{2}\leq x\leq 1,\end{cases}$
and
$h_{1}(x)\coloneqq\frac{3}{16}x+\frac{13}{16},\qquad 0\leq x\leq 1.$
Then the functions $h_{0}$ and $h_{1}$ are strictly increasing and continuous. Moreover,
$0=h_{0}(0)<h_{0}(1)=\frac{13}{16}=h_{1}(0)<h_{1}(1)=1.$
Hence, $\{h_{0},h_{1}\}$ is a compatible system.
We also observe that
$h_{0}(x)\geq\frac{3}{8}\quad\text{for every }x\in\left[\frac{3}{8},1\right].$
Therefore,
$\lim_{n\to\infty}h_{0}^{n}(1)\geq\frac{3}{8}.$
Since $\displaystyle\inf\mathcal{D}_{h}\setminus\{0\}=\lim_{n\to\infty}h_{0}^{n}(1)$, it follows that $\mathcal{D}_{h}\cap(0,\frac{3}{8})=\emptyset$. Consequently, $\mathcal{D}_{h}$ is not dense in $[0,1]$.
Finally, we remark that $h_{0}$ has a fixed point other than zero; indeed, $\displaystyle h_{0}\!\left(\frac{3}{8}\right)=\frac{3}{8}$.
It is natural to consider the inverse function of the solution $\varphi$. We easily see the following assertion.
@@
Assume that $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are D-systems on $[0,1]$. Let $\varphi$ be the unique solution of (1.1). Then, the inverse function $\varphi^{-1}$ of $\varphi$ is well-defined and satisfies the following conjugate equation
@@
which is equivalent with (1.1) where $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are exchanged.
By [6, Proposition 3.1], if $\varphi$ is singular, then $\varphi^{-1}$ is also singular.